OpenStudy (anonymous):

How many moles of gold are in 2.23 X 10^17 atoms of gold?

OpenStudy (anonymous):

@nadaddy Do you know how many atoms are in a mole?

OpenStudy (anonymous):

not really im sort of new with this subject

OpenStudy (anonymous):

No problem. A mole is just a quantity (just like a dozen is a quantity). The mole is equal to \[6.022 \times 10^{22}\] So a mole of x will have 6.022 x 10^22 of x. 2 mole of x will have 12.044 x 10^22 of x 0.5 mole of x will have 3.011 x 10^22 of x A mole of atoms will have 6.022 x 10^22 atoms

OpenStudy (anonymous):

You were given 2.23 x 10^27 atoms so you want to convert atoms to moles. If you set up your problem so that atoms and atoms cancel you will be left with moles. Meaning... \[\frac{ 2.23 \times 10^{27}atoms }{ 1 }\times \frac{1 mole }{ atoms }\] atoms in the numerator and atoms in the denominator will cancel and you'll be left with moles. You just need to fill in the correct number by atoms and calculate it.

OpenStudy (anonymous):

so how do we determine the number of atoms and do we plug that number into the fraction and multiply it by 2.23 x 10^27?

OpenStudy (anonymous):

You are trying to solve the number of moles. 2.23 x 10^27 is in the numerator. The number of atoms per mole goes in the denominator.

OpenStudy (anonymous):

im sorry im uneducated when it comes to this stuff chemistry is really tough for me :( but in order to find the number of moles do we need find the number of atoms or is that already given to us?

OpenStudy (anonymous):

Your question states "How many moles of gold are in 2.23 X 10^17 atoms of gold? " The number of atoms is given to you "2.23 atoms."

OpenStudy (anonymous):

My mistake "2.23 x 10^27 atoms"

OpenStudy (anonymous):

so going back to the fraction, how do we fill in the correct number by atoms and calculate it?

OpenStudy (anonymous):

In one mole there is 6.022 x 10^23 atoms, this number goes in the denominator where I wrote atoms but didn't put a number.

OpenStudy (anonymous):

so we would plug 6.022 x 10

OpenStudy (anonymous):

wait i accidentally hit post hold on one sec

OpenStudy (anonymous):

so we would plug 6.022 x 10 into the numerator of the second fraction correct?

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

and once we have all the numbers plugged into the fraction, the product would be our answer?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So if the atoms cancel each other out am i left with 6.022 x 10^23/1?

OpenStudy (anonymous):

No. Another wat of writing what I wrote earlier is \[\frac{ 2.23 \times 10^{17} atoms \times mole }{6.022 \times 10^{23} atoms }\] Atoms and atoms cancel meaning the unit atom cancel each othe out but not the actual numbers